# Measuring the Substructure Mass Power Spectrum of 23 SLACS Strong Galaxy-Galaxy Lenses with Convolutional Neural Networks

### v1

@@ 4.2 Convolutional neural network architecture
The CNN architecture used in Vernardos et al. (2020) consisted of a series of 7 convolutional layers that were then flattened and followed by 2 fully connected layers for each parameter. In this work we use a variant of ResNet-18 (He et al., 2015) to include residual skip connections into our CNN. The addition of skip connections enables CNNs to be significantly deeper without suffering from vanishing gradient and degradation issues and has been shown to improve gradient flow and training stability. We also use squeeze and excitation connections (Hu et al., 2017) to add a spatial attention mechanism to our network. The input of the CNN is $110\times 110$ pixel images ($5.5^{\prime\prime}$ across with a resolution of $0.05^{\prime\prime}$ per pixel to match the SLACS data). The standard ResNet-18 starts with one convolutional layer with a filter size of 64 and 4 residual blocks with 4 convolutional layers each and filter sizes of [64, 128, 256, 512] respectively. We use the same architecture except with smaller filter sizes of [32, 64, 128, 256] to use fewer parameters and avoid overfitting and the addition of the squeeze and excitation connections with a ratio of eight (He et al., 2015; Hu et al., 2017). The output of the final convolutional layer is flattened and split into two separate fully connected layers with $N_{p}=100$ classes, representing the bins of the discrete probability distributions of our output in $\log_{10}(\sigma_{\delta\psi}^{2})$ and $\beta$. The final fully connected layers are deterministic instead of Bayesian since this has been shown to improve performance (Hortúa et al., 2020). We also includes a softmax activation function to normalize the final outputs to a probability distributions such that all classes are positive and sum to one. Our model [^4] https://github.com/JFagin/GRF_ML has 5,735,456 trainable parameters that are initialized randomly and is implemented using TensorFlow and the TensorFlow-Probability API (Abadi et al., 2015).
@@ 4.3 Training and performance
In Vernardos et al. (2020) they used a value of $\lambda=0.9$, but we use a value of $\lambda=0.975$, which leads to the entropy term contributing $\approx 45$ per cent of the relative contribution to the loss. Our value for $\lambda$ is purposely closer to 1 in order to avoid learning narrower predicted probability distributions and thus underestimating the uncertainty. In addition to equation (7), the regularization term of the BNN described in Sect. 4.1.2 is appended to the loss to control the standard deviation of the model weights.
@@
For each data set, we allocate 230,000 of the mock images for training the CNN while reserving 20,000 for validation. We train the CNNs through multiple passes of the entire training sets, referred to as epochs. Each CNN is trained for a total of 2,000 epochs. During both training and inference, we augment each image by randomly rotating in intervals of $90\degree$ and flipping vertically or horizontally. This teaches the CNNs to remain invariant under these transformations and effectively increases the number of training examples by a factor of 8, preventing overfitting. We find that augmenting the data during inference improves uncertainty estimates as we average the predictions across various orientations. We also sample a new realization of the training labels each epoch, as mentioned in Sect. 4.1.1.
@@
Our CNNs are trained by minimizing the loss function using an Adam optimizer (Kingma & Ba, 2017) with an initial learning rate of $10^{-3}$ and a batch size of 250. The learning rate is exponentially decayed to $10^{-4}$ over the course of training. Figure 9 shows how the mean squared error (MSE) between the mean of each predicted probability distribution and the ground truth varies with epoch across the validation set of the combined data set. We note that the MSE is indirectly minimized because the CNNs are never shown the ground truth values directly. The parameter space is scaled from zero to one for comparison, making the root MSE effectively represent a percentage error relative to the total parameter space. The validation MSE levels off by 2,000 epochs, so further training could lead to overfitting and be detrimental to the model performance. After the CNNs are trained, we apply them to the validation set to evaluate their performance. We report a final root MSE of 20.9 per cent for $\log_{10}(\sigma_{\delta\psi}^{2})$ and 20.1 per cent for $\beta$ using the combined data set. Additional metrics and figures for data set 1 and 2 are given in Appendix B and show very similar performance.
@@
Figure 10: Mean of our predicted probability distributions for $\log_{10}(\sigma_{\delta\psi}^{2})$ (top) and $\beta$ (bottom) compared to the ground truth across the validation set of the combined data set. The diagonal blue line represents the ideal case where each mean prediction would match the true value.
@@
Figure 11: Histogram of the difference between the mean of our predicted probability distributions and the ground truth across the validation set of the combined data set. The ideal case where the mean prediction matches the ground truth occurs at the centre, i.e. when $\Delta\,\log_{10}(\sigma_{\delta\psi}^{2})$ and $\Delta\,\beta$ are zero.
@@
Figure 12: Coverage probabilities for the combined data set, i.e. the fraction of the validation set which has ground truth values within 68, 95, and 99 per cent confidence intervals. The ideal case of perfect uncertainty calibration, when the coverage probabilities exactly match the middle confidence intervals, is shown by the horizontal dashed lines.
@@
The coverage probabilities are the fraction of ground truth values that lie within a given confidence interval of our predicted probability distributions across the validation set. To assess the validity of our predicted probability distributions as uncertainties, Fig. 12 shows the coverage probabilities for the 68, 95, and 99 per cent confidence intervals. We assess these confidence intervals since they represent the $1\sigma$, $2\sigma$, and $3\sigma$ confidence intervals of a Gaussian, although our probability distributions can take on any form. Ideally, all the ground truth values would be within our reported confidence intervals at the same rate as our predictions. The converge probabilities nearly line up with the ideal case, so we can conclude that our predicted probability distributions approximately represent the uncertainty in our predictions. For better calibration of our uncertainties, we set $n=0.9$ in equation (6) and $\lambda=0.975$ in equation (7).
@@ 6 Discussion and conclusions
We demonstrated how ML can be used to predict the power-law statistics of a GRF perturbation to the lens potential by expanding on the framework of Vernardos et al. (2020) to make it more viable and reliable for observed data. In this study we used HST images of 23 SLACS lenses, which will be similar to Euclid in terms of data quality and resolution. To apply ML to the tens of thousands of lenses expected to be found by Euclid and the Vera C. Rubin observatory, a sufficiently diverse training set should be produced. For example, sources could be reconstructed by other ML methods such as Biggio et al. (2022); Karchev et al. (2022); Adam et al. (2023) to build a training set that encapsulates all the known lenses. This could be combined with real galaxy images, simple analytic brightness profiles, or galaxy images from generative deep learning models (e.g. Lanusse et al., 2021). Ideally multiple training sets would be produced with different sources to ensure that the resulting predictions are independent of any particular assumption, like the training sets constructed in this work. A larger population of lenses will allow us to marginalize over the uncertainty in our predictions and more tightly constrain the joint probability distribution of the GRF power-law parameters, i.e. by narrowing the predictions in Fig. 15.
In this study, we assumed an SIE plus external shear smooth lens model and a GRF perturbation parameterized by a power-law, but our method could be used to quantify any type of perturbation on a smooth lens mass model. Consistent with Chatterjee & Koopmans (2017); Vernardos et al. (2020); Bayer et al. (2023a, b), we parameterized the GRF power spectrum using a simple power-law (see equation (2)), but a better parameterization could potentially be developed by analysing the power spectrum of numerical N-body and hydrodynamic simulations with different dark matter models. We also note that the GRF statistical framework works well at approximating the collective contribution of smaller mass dark matter substructure, but if the lens perturbation is dominated by very few large mass dark matter subhaloes, then this approximation breaks down (Vegetti et al., 2010; Hezaveh et al., 2016b; Chatterjee & Koopmans, 2017). In future work, the low mass dark matter subhaloes could be modelled as a GRF but include possible contamination by the occasional large mass subhalo that can have a substantial localized effect on the lens potential. Improvements could also be made to disentangle baryonic substructure from the dark matter substructure with the inclusion of boxyness, discyness, or other baryonic processes in the smooth mass profile (Van de Vyvere et al., 2022).
@@
While we showed that our reported uncertainty range is appropriate across the validation set of mock images, there may be additional systematic uncertainties when applying the CNN to real lenses which we cannot fully account for. We partially addressed this by training three separate models using different source methods, but other choices in building the training set could affect our results. For example, when simulating our mock lenses we assumed a Sérsic lens light profile, while the real data may have more complicated lens light profiles which often require 2 Sérsic profiles to fit. Discrepancies in uncertainty estimations may always exist when a neural network is trained and validated with a simulated data set and then applied to real data. In this work we attempted to bridge some of these discrepancies by using separately trained CNNs with complementary training sets containing different sources, which Vernardos et al. (2020) found to be the most important consideration. Further diversification could also be used by altering the smooth lens model, the lens light model, or the observational effects across training sets.
@@
For future observations, a combination of ML approaches like ours and more traditional lens modelling approaches like in Bayer et al. (2023b) may give better results than using either method individually. ML could be used to first give a quick estimate of the GRF perturbation parameters across the tens of thousands of lenses expected to be observed by next generation surveys. Lenses with interesting properties could be selected for follow up studies using a computationally expensive lens modelling approach to further constrain the lens properties and double-check the ML predictions. As of now however, the method of (Bayer et al., 2023b) only produces upper bound constraints on the GRF parameters while our ML method can make direct predictions. In future work, our predictions could also be compared with the predictions from hydrodynamic simulations such as EAGLE (Schaye et al., 2014) and IllustrisTNG (Nelson et al., 2021).
@@ Appendix B Additional figures from different data sets
Figure 18 shows additional figures from the CNNs trained on data set 1 and 2. This includes the MSE as a function of epoch, the mean predictions compared to the ground truth, difference histograms between the mean predictions and the ground truth, and the coverage probabilities for the 68, 95, and 99 per cent confidence intervals. Figure 19 gives the average predicted probability distributions across our three CNNs for each of the SLACS lenses. Figures 2021, and 22 give our predictions compared to the smooth lensing parameters for our CNNs trained on data set 1, data set 2, and the combined data set respectively. Data set 1 (data set 2) has a root MSE of 20.8 (20.9) per cent for $\log_{10}(\sigma_{\delta\psi}^{2})$ and 19.7 (20.0) per cent for $\beta$.

### v2

@@ 4.2 Convolutional neural network architecture
The CNN architecture used in Vernardos et al. (2020) consisted of a series of 7 convolutional layers that were then flattened and followed by 2 fully connected layers for each parameter. In this work we use a variant of ResNet-18 (He et al., 2015) to include residual skip connections into our CNN. The addition of skip connections enables CNNs to be significantly deeper without suffering from vanishing gradient and degradation issues and has been shown to improve gradient flow and training stability. We also use squeeze and excitation connections (Hu et al., 2017) to add a spatial attention mechanism to our network. The input of the CNN is $110\times 110$ pixel images ($5.5^{\prime\prime}$ across with a resolution of $0.05^{\prime\prime}$ per pixel to match the SLACS data). The standard ResNet-18 starts with one convolutional layer with a filter size of 64 and 4 residual blocks with 4 convolutional layers each and filter sizes of [64, 128, 256, 512] respectively. We use the same architecture except with smaller filter sizes of [32, 64, 128, 256] to use fewer parameters and avoid overfitting, and we include the addition of the squeeze and excitation connections with a ratio of eight (He et al., 2015; Hu et al., 2017). The output of the final convolutional layer is flattened and split into two separate fully connected layers with $N_{p}=100$ classes, representing the bins of the discrete probability distributions of our output in $\log_{10}(\sigma_{\delta\psi}^{2})$ and $\beta$. The final fully connected layers are deterministic instead of Bayesian since this has been shown to improve performance (Hortúa et al., 2020). We also includes a softmax activation function to normalize the final outputs to a probability distributions such that all classes are positive and sum to one. Our model [^4] https://github.com/JFagin/GRF_ML has 5,735,456 trainable parameters that are initialized randomly and is implemented using TensorFlow and the TensorFlow-Probability API (Abadi et al., 2015).
@@ 4.3 Training and performance
In Vernardos et al. (2020) they used a value of $\lambda=0.9$, but we use a value of $\lambda=0.975$, which leads to the entropy term contributing $\approx 45$ per cent of the relative contribution to the loss. Our value for $\lambda$ is purposely closer to 1 in order to avoid learning narrower predicted probability distributions and thus underestimating the uncertainty. In addition to equation (7), the regularization term of the BNN described in Section 4.1.2 is appended to the loss to control the standard deviation of the model weights.
@@
We train the CNNs through multiple passes of the entire training sets, referred to as epochs. Each CNN is trained for a total of 2,000 epochs. During both training and inference, we augment each image by randomly rotating in intervals of $90\degree$ and flipping vertically or horizontally. This teaches the CNNs to remain invariant under these transformations and effectively increases the number of training examples by a factor of 8, preventing overfitting. We find that augmenting the data during inference improves uncertainty estimates as we average the predictions across various orientations. We also sample a new realization of the training labels each epoch, as mentioned in Section 4.1.1.
@@
Our CNNs are trained by minimizing the loss function using an Adam optimizer (Kingma & Ba, 2017) with an initial learning rate of $10^{-3}$ and a batch size of 250. The learning rate is exponentially decayed to $10^{-4}$ over the course of training. Figure 9 shows how the mean squared error (MSE) between the mean of each predicted probability distribution and the ground truth varies with epoch across the validation set of the combined data set. We note that the MSE is indirectly minimized because the CNNs are never shown the ground truth values directly. The parameter space is scaled from zero to one for comparison, making the root MSE effectively represent a percentage error relative to the total parameter space. The validation MSE levels off by 2,000 epochs, so further training could lead to overfitting and be detrimental to the model performance. After the CNNs are trained, we apply them to the test sets to evaluate their performance. We report a final root MSE of 20.6 per cent for $\log_{10}(\sigma_{\delta\psi}^{2})$ and 19.9 per cent for $\beta$ using the combined data set. We estimate the relative contribution of the epistemic and aleatoric uncertainties by comparing the standard deviation of our probability distributions for a single prediction and the average of 200 predictions for each test set example. We find that the standard deviation increased by 9.2 per cent for $\log_{10}(\sigma_{\delta\psi}^{2})$ and 5.4 per cent for $\beta$ for the combined data set, which represents the percentage contribution of the epistemic uncertainty as compared to the aleatoric uncertainty. We expect the contribution of the epistemic uncertainty to be small compared to the aleatoric uncertainty since it should decrease with the size of the training set, and we use a relatively large training set of 220,000 images. Additional metrics and figures for data set 1 and 2 are given in Appendix B and show similar performance.
@@
Figure 10: Mean of our predicted probability distributions for $\log_{10}(\sigma_{\delta\psi}^{2})$ (top) and $\beta$ (bottom) compared to the ground truth across the test set of the combined data set. The diagonal blue line represents the ideal case where each mean prediction would match the true value.
@@
Figure 11: Histogram of the difference between the mean of our predicted probability distributions and the ground truth across the test set of the combined data set. The ideal case where the mean prediction matches the ground truth occurs at the centre, i.e. when $\Delta\,\log_{10}(\sigma_{\delta\psi}^{2})$ and $\Delta\,\beta$ are zero.
@@
Figure 12: Coverage probabilities for the combined data set, i.e. the fraction of the test set which has ground truth values within 68, 95, and 99 per cent confidence intervals. The ideal case of perfect uncertainty calibration, when the coverage probabilities exactly match the middle confidence intervals, is shown by the horizontal dashed lines.
@@
The coverage probabilities are the fraction of ground truth values that lie within a given confidence interval of our predicted probability distributions across the test set. To assess the validity of our predicted probability distributions as uncertainties, Fig. 12 shows the coverage probabilities for the 68, 95, and 99 per cent confidence intervals. We assess these confidence intervals since they represent the $1\sigma$, $2\sigma$, and $3\sigma$ confidence intervals of a Gaussian, although our probability distributions can take on any form. Ideally, all the ground truth values would be within our reported confidence intervals at the same rate as our predictions. The converge probabilities nearly line up with the ideal case, so we can conclude that our predicted probability distributions approximately represent the uncertainty in our predictions. For better calibration of our uncertainties, we set $n=0.9$ in equation (6) and $\lambda=0.975$ in equation (7).
@@ 6 Discussion and conclusions
We demonstrated how ML can be used to predict the power-law statistics of a GRF perturbation to the lens potential by expanding on the framework of Vernardos et al. (2020) to make it more viable and reliable for observed data. In this study we used HST images of 23 SLACS lenses, which will be similar to Euclid in terms of data quality and resolution. To apply ML to the tens of thousands of lenses expected to be found by Euclid and the Vera C. Rubin observatory, a sufficiently diverse training set should be produced. For example, sources could be reconstructed by other ML methods such as Biggio et al. (2022); Karchev et al. (2022); Adam et al. (2023) to build a training set that encapsulates all the known lenses. This could be combined with real galaxy images, simple analytic brightness profiles, or galaxy images from generative deep learning models (e.g., Lanusse et al., 2021). Ideally multiple training sets would be produced with different sources to ensure that the resulting predictions are independent of any particular assumption, like the training sets constructed in this work. A larger population of lenses will allow us to marginalize over the uncertainty in our predictions and more tightly constrain the joint probability distribution of the GRF power-law parameters, i.e. by narrowing the predictions in Fig. 15.
In this study, we assumed an SIE plus external shear smooth lens model and a GRF perturbation parameterized by a power-law, but our method could be used to quantify any type of perturbation on a smooth lens mass model. Consistent with Chatterjee & Koopmans (2017); Vernardos et al. (2020); Bayer et al. (2023a, b), we parameterized the GRF power spectrum using a simple power-law (see equation (2)), but a better parameterization could potentially be developed by analysing the power spectrum of numerical N-body and hydrodynamic simulations with different dark matter models. We also note that the GRF statistical framework works well at approximating the collective contribution of smaller mass dark matter substructure, but if the lens perturbation is dominated by very few large mass dark matter subhaloes, then this approximation breaks down (Vegetti et al., 2010; Hezaveh et al., 2016b; Chatterjee & Koopmans, 2017). In future work, the low mass dark matter subhaloes could be modelled as a GRF but include possible contamination by the occasional large mass subhalo that can have a substantial localized effect on the lens potential. Improvements could also be made to disentangle baryonic substructure from the dark matter substructure with the inclusion of boxyness, discyness, or other baryonic processes in the smooth mass profile (see, Van de Vyvere et al., 2022).
@@
While we showed that our reported uncertainty range is appropriate across the test set of mock images, there may be additional systematic uncertainties when applying the CNN to real lenses which we cannot fully account for. We partially addressed this by training three separate models using different source methods, which Vernardos et al. (2020) found to be the most important consideration. The sources from Shajib et al. (2021) may deviate from the true sources of our lenses. In data set 2 we used the entire MCMC chain, so any source discrepancies would have to be completely outside the posterior distribution of the MCMC chain to bias our network. We also compare the results of using the network trained on the entire MCMC chain to the network trained with sources produced through data augmentation of the best fitting sources and find agreement in the results. The data augmentation procedure can lead to large deviations in the source morphology, so the true sources would have to be completely outside that posterior distribution as well. Other choices in building the training set could also affect our results. For example, when simulating our mock lenses we assumed a Sérsic lens light profile, while the real data may have more complicated lens light profiles which often require 2 Sérsic profiles to fit. Discrepancies in uncertainty estimations may always exist when a neural network is trained and tested with simulated data and then applied to real data. We attempted to bridge some of these discrepancies by using separately trained CNNs with complementary training sets containing different sources and then averaging the resulting predicted probability distributions, increasing the reported uncertainty. Further diversification could also be used by altering the smooth lens model, the lens light model, or the observational effects across training sets.
@@
For future observations, a combination of ML approaches like ours and more traditional lens modelling approaches like in Bayer et al. (2023b) may give better results than using either method individually. A pretrained neural network could be used to first give an estimate of the GRF perturbation parameters across the tens of thousands of lenses expected to be observed by next generation surveys in just a few minutes of computational time with a single GPU. Lenses with interesting properties could be selected for follow up studies using a computationally expensive lens modelling approach to further constrain the lens properties and double-check the ML predictions. These methods can take weeks of computational time and often need to be fine tuned to each system, making them unrealistic to apply to all the lenses. As of now, the method of (Bayer et al., 2023b) only produces upper bound constraints on the GRF parameters while our ML method can make direct predictions. In future work, our predictions could also be compared with the predictions from hydrodynamic simulations such as EAGLE (Schaye et al., 2014) and IllustrisTNG (Nelson et al., 2021).
@@ Appendix B Additional figures from different data sets
Figure 18 shows additional figures from the CNNs trained on data set 1 and 2. This includes the MSE as a function of epoch, the mean predictions compared to the ground truth, difference histograms between the mean predictions and the ground truth, and the coverage probabilities for the 68, 95, and 99 per cent confidence intervals. Figure 19 gives the average predicted probability distributions across our three CNNs for each of the SLACS lenses. Figures 2021, and 22 give our predictions compared to the smooth lensing parameters measured by Shajib et al. (2021) for our CNNs trained on data set 1, data set 2, and the combined data set respectively. Data set 1 (data set 2) has a root MSE of 20.1 (20.7) per cent for $\log_{10}(\sigma_{\delta\psi}^{2})$ and 19.5 (20.1) per cent for $\beta$. The relative contribution of the epistemic uncertainties compared to the aleatoric uncertainty for data set 1 (data set 2) is 9.0 (8.9) per cent for $\log_{10}(\sigma_{\delta\psi}^{2})$ and 5.5 (5.6) per cent for $\beta$.